# Differentiable Pareto-Smoothed Weighting for High-Dimensional Heterogeneous Treatment Effect Estimation

### v1

@@ 5 Related Work
Data-driven feature separation for CATE estimation: CATE estimation has gained increasing attention because of its great importance for causal mechanism understanding [Chikahara et al., 2022, Zhao et al., 2022] and for decision support in various fields, such as precision medicine [Gao et al., 2021] and online advertising [Sun et al., 2015]. There has been a surge of interest in leveraging flexible machine learning models, including tree-based models [Athey et al., 2019, Hill, 2011], Gaussian processes [Alaa and van der Schaar, 2018, Horii and Chikahara, 2024], and neural networks [Hassanpour and Greiner, 2019, Johansson et al., 2016, Shalit et al., 2017]. However, most methods treat all input features X as confounders. As pointed out by Wu et al. [2022], the empirical performance of such methods varies a lot with the presence of adjustment variables in X, which is usual in practice, especially in high-dimensional settings.
@@
This issue motivates us to develop data-driven feature separation methods for treatment effect estimation. A pioneer work is the data-driven variable decomposition (D2VD) [Kuang et al., 2017, 2020], which minimizes the weighted prediction loss plus the regularizer for feature separation. The recent method addresses a more complicated setup, where features X include post-treatment variables, which are affected by treatment $A$ [Wang et al., 2023]. However, the estimation target of these methods is ATE, not CATE.
@@
By contrast, DRCFR deals with CATE estimation and is founded on weighted representation learning, which is a promising approach for addressing high-dimensional data. This is why we adopted it as the inference engine of our weight correction framework. Integrating the recent idea of enforcing independence between feature representations with mutual information [Cheng et al., 2022, Chu et al., 2022, Liu et al., 2024] is left as our future work.
@@
Weighting schemes for treatment effect estimation: IPW [Rosenbaum and Rubin, 1983] is a common weighting technique for treatment effect estimation. However, the weighted estimator based on IPW is often numerically unstable due to the computation of the inverse of the propensity score. One remedy for this issue is weight truncation [Crump et al., 2009], which, however, causes the estimation bias, leading to inaccurate treatment effect estimation.
@@
To improve the estimation performance, Zhu et al. [2020] have employed Pareto smoothing [Vehtari et al., 2024]. Although they empirically show that using this technique leads to better performance than weight truncation, their method is developed for the estimation of ATE, not CATE.
@@
Applying Pareto smoothing in weighted representation learning for CATE estimation is difficult because it prevents gradient backpropagation due to the non-differentiability. This difficulty is disappointing, given that previous work has theoretically shown that the weight correction schemes, such as weight truncation, helps to extract predictive feature representations for CATE estimation [Assaad et al., 2021].
@@
To establish a Pareto-smoothed weighting framework for CATE estimation from high-dimensional data, we have shown that how differentiable ranking technique [Blondel et al., 2020] can be used to simultaneously learn the propensity score model and the feature representations.
@@ 6 Conclusion
In this paper, we have established a differentiable Pareto-smoothed weighting framework for CATE estimation from high-dimensional data. To construct a CATE estimator that is numerically robust to propensity score estimation error, we make the weight correction procedure in Pareto smoothing differentiable and incorporate it into the weighted representation learning approach for CATE estimation. Experimental results show that our framework successfully outperforms traditional weighting schemes, as well as the existing CATE estimation methods.
Leveraging the versatility of weighting, our future work will investigate how to extend our framework to estimate the effects of such complex treatment as high-dimensional binary treatment [Zou et al., 2020], continuous-valued treatment [Wang et al., 2022], time series treatment [Lim et al., 2018].
@@ Appendix A Rank Function Definition
In Section 3.3, we consider rank function $r$, which takes input vector $\textbf{{w}}=[w_{1},\dots,w_{n}]^{\top}$ and outputs the rank of each element in w. We formally define this rank function based on the sorting operation and the concept of inverse permutation.
Consider a sorting operation over $w_{1},\dots,w_{n}$ in $\textbf{{w}}\in\operatorname{\mathbb{R}}^{n}$ that finds a permutation $\operatorname{\bm{\rho}}=[\rho_{1},\dots,\rho_{n}]^{\top}$ such that the vector values permuted according to $\operatorname{\bm{\rho}}$, $\textbf{{w}}_{\operatorname{\bm{\rho}}}=[w_{\rho_{1}},\dots,w_{\rho_{n}}]^{\top}$, are increasing as $w_{\rho_{1}}\leq\dots\leq w_{\rho_{n}}$. Let $\operatorname{\bm{\rho}}^{-1}$ be the inverse of permutation $\operatorname{\bm{\rho}}$, i.e., a permutation whose $\rho_{i}$-th element is $i$ for $i=1,\dots,n$.
@@
Throughout this paper, we consider rank function $r(\textbf{{w}})$ that evaluates the position of each $w_{i}$ based on the sorting in an ascending order. If we need to consider the ranking in descending order, we can formulate it as $r(-\textbf{{w}})$.
@@ B.1 Settings of Baselines
Regarding linear regression methods (LR-1 and LR-2), to avoid overfitting due to the large number of input features, we employed the ridge regression model in scikit-learn [Pedregosa et al., 2011]. For the meta-learner methods (i.e., the S-Learner (SL), the T-Learner (TL), the X-Learner (XL), the DR-Learner (DRL)) and the tree-based methods (CF and CF DML) we used the EconML Python package [Battocchi et al., 2019]. As the base learners for SL, TL, XL, and DRL, we chose random forest because we empirically observed that it achieved the best performance among the three model candidates, random forest, gradient boosting, and support vector machine.
We evaluated the performance of neural network methods (TARNet and GANITE), using the existing implementations [Curth and van der Schaar, 2021, Yoon et al., 2018].  [^4] https://github.com/AliciaCurth/CATENets  [^5] https://github.com/jsyoon0823/GANITE As regards the DRCFR method, we implemented it using PyTorch [Paszke et al., 2019].
With PSW, we trained propensity score model $\pi(\textbf{{X}})$ beforehand, using a paired sample {$(a_{i},\textbf{{x}}_{i})$} $\subset\mathcal{D}$. Here we formulated $\pi(\textbf{{X}})$ using the three-layered FNN, as with our method and DRCFR. After having computed the IPW weights based on the trained propensity score model, we performed Pareto smoothing over them, using psislw() function in the Python package named ArviZ [Kumar et al., 2019]. Using the Pareto-smoothed weights, we learned the two encoders $\Delta(\textbf{{X}})$ and $\Upsilon(\textbf{{X}})$, as well as outcome prediction models $h^{0}$ and $h^{1}$.
@@ B.2 Synthetic Data
Following Hassanpour and Greiner [2020], we prepared the synthetic datasets.  [^6] We used their implementation in https://www.dropbox.com/sh/vrux2exqwc9uh7k/AAAR4tlJLScPlkmPruvbrTJQa?dl=0.
We first drawn the values of features $\textbf{{X}}\in\operatorname{\mathbb{R}}^{d}$ from standard multivariate Gaussian distribution:
@@
Then, we sampled the values of treatment $A$ and outcome $Y$ using $\textbf{{X}}_{\Gamma}\in\operatorname{\mathbb{R}}^{d/3}$, $\textbf{{X}}_{\Delta}\in\operatorname{\mathbb{R}}^{d/3}$, and $\textbf{{X}}_{\Upsilon}\in\operatorname{\mathbb{R}}^{d/3}$, which are the feature subsets in $\textbf{{X}}=[\textbf{{X}}_{\Gamma},\textbf{{X}}_{\Delta},\textbf{{X}}_{\Upsilon}]^{\top}$. In particular, we employed $\psi=[\textbf{{X}}_{\Gamma},\textbf{{X}}_{\Delta}]^{\top}\in\operatorname{\mathbb{R}}^{2d/3}$, to generate $A$’s values as
@@
where $\textbf{{c}}_{Y^{0}}\in\operatorname{\mathbb{R}}^{2d/3}$ and $\textbf{{c}}_{Y^{1}}\in\operatorname{\mathbb{R}}^{2d/3}$ are the coefficient vectors drawn from $\mathcal{N}(0,1)$, symbol $\odot$ denotes element-wise product (a.k.a., Hadamard product), and $\epsilon\sim\mathcal{N}(0,1)$ is a scalar standard Gaussian noise. Using the values of potential outcomes $Y^{0}$ and $Y^{1}$, we computed the values of outcome $Y$ by $Y=AY^{1}+(1-A)Y^{0}$.
@@ Appendix C Additional Experimental Results
In Section 4.2, as with Hassanpour and Greiner [2020], we present the performance on the relatively low-dimensional synthetic data. To further investigate the performance of our method, we performed additional experiments using high-dimensional synthetic data.
Here, we generated the synthetic data in the same way (as described in Section B.2) except that the number of feature $d$ was set to $d=600,1200,\dots,3000$.
Figure 4 shows the results. Again, our method achieved better CATE estimation performance than DRCFR and TARNet, thus showing that it can successfully improve the estimation performance of DRCFR by effectively correcting the IPW weights via differentiable Pareto smoothing.
Regarding the feature representation learning performance (Figure 4 (a) - (c)), both DRCFR and our method worked better than TARNet. As illustrated in Figure 4 (a), however, with DRCFR and our method, the difference among the absolute parameter values in learned encoder $\Gamma(\textbf{{X}})$ decreased, as the number of features $d$ increased, indicating that distinguishing the representation of instrumental variables $\Gamma(\textbf{{X}})$ from the one of confounders $\Delta(\textbf{{X}})$ was difficult for both methods. This difficulty is partly because binary treatment $A$ can be accurately predicted solely from the representation of confounders (i.e., $\Delta(\textbf{{X}})$) when there are a sufficient number of confounders $\textbf{{X}}_{\Delta}$.
@@
Figure 4: Learned encoder parameter differences and test PEHEs on synthetic data. (a): Value difference of $\mathbf{W}^{1}$ in encoder $\Gamma(\textbf{{X}})$; (b): Value difference of $\mathbf{W}^{1}$ in encoder $\Delta(\textbf{{X}})$; (c): Value difference of $\mathbf{W}^{1}$ in encoder $\Upsilon(\textbf{{X}})$; (d) Test PEHEs. With TARNet, since it learns single encoder, we computed all parameter value differences with weight matrix in same encoder.

### v2

@@ 5 Related Work
Data-driven feature separation for CATE estimation: CATE estimation has gained increasing attention because of its great importance for causal mechanism understanding [Chikahara et al., 2022, Zhao et al., 2022] and for decision support in various fields, such as precision medicine [Gao et al., 2021] and online advertising [Sun et al., 2015]. There has been a surge of interest in leveraging flexible machine learning models, including tree-based models [Athey et al., 2019, Hill, 2011], Gaussian processes [Alaa and van der Schaar, 2018, Horii and Chikahara, 2024], and neural networks [Hassanpour and Greiner, 2019, Johansson et al., 2016, Shalit et al., 2017]. However, most methods treat all input features X as confounders. As pointed out by Wu et al. [2022], the empirical performance of such methods greatly varies with the presence of adjustment variables in X, which is usual in practice, especially in high-dimensional settings.
@@
Motivated by this issue, we develop data-driven feature separation methods for treatment effect estimation. A pioneering work is data-driven variable decomposition (D2VD) [Kuang et al., 2017, 2020], which minimizes the weighted prediction loss plus the regularizer for feature separation. A recent method addresses a more complicated setup, where features X include post-treatment variables, which are affected by treatment $A$ [Wang et al., 2023]. However, the estimation target of these methods is ATE, not CATE.
@@
By contrast, DRCFR deals with CATE estimation and is founded on weighted representation learning, which is a promising approach for addressing high-dimensional data. These advantages are why we adopted it as the inference engine of our weight correction framework. Integrating the recent idea of enforcing independence between feature representations with mutual information [Cheng et al., 2022, Chu et al., 2022, Liu et al., 2024] remains our future work.
@@
Weighting schemes for treatment effect estimation: IPW [Rosenbaum and Rubin, 1983] is a common weighting technique for treatment effect estimation. However, a weighted estimator based on IPW is often numerically unstable due to the computation of the inverse of propensity scores. One remedy is weight truncation [Crump et al., 2009], which, however, causes estimation bias, leading to inaccurate treatment effect estimation.
@@
To improve the estimation performance, Zhu et al. [2020] employed Pareto smoothing [Vehtari et al., 2024]. Although they empirically show that using this technique leads to better performance than weight truncation, their method was developed to estimate ATE, not CATE.
@@
Applying Pareto smoothing in weighted representation learning for CATE estimation is difficult because it prevents gradient backpropagation due to non-differentiability. This difficulty is disappointing, given that previous work has theoretically shown that such weight correction schemes as weight truncation help extract predictive feature representations for CATE estimation [Assaad et al., 2021].
@@
To establish a Pareto-smoothed weighting framework for CATE estimation from high-dimensional data, we demonstrated how a differentiable ranking technique [Blondel et al., 2020] can be used to simultaneously learn a propensity score model and feature representations.
@@ 6 Conclusion
We established a differentiable Pareto-smoothed weighting framework for CATE estimation from high-dimensional data. To construct a CATE estimator that is numerically robust to propensity score estimation error, we develop a differentiable weight correction procedure based on Pareto smoothing and incorporated it into weighted representation learning for CATE estimation. We experimentally show that our framework outperformed traditional weighting schemes as well as the existing CATE estimation methods.
By leveraging the versatility of weighting, our future work will investigate how to extend our framework to estimate the effects of high-dimensional binary treatment [Zou et al., 2020], continuous-valued treatment [Wang et al., 2022], and time series treatment [Lim et al., 2018].
@@ Appendix A Rank Function Definition
In Section 3.3, we consider rank function $r$, which takes input vector $\textbf{{w}}=[w_{1},\dots,w_{n}]^{\top}$ and outputs the rank of each element in w. We formally define this rank function based on a sorting operation and the concept of inverse permutation.
Consider a sorting operation over $w_{1},\dots,w_{n}$ in $\textbf{{w}}\in\operatorname{\mathbb{R}}^{n}$ that finds permutation $\operatorname{\bm{\rho}}=[\rho_{1},\dots,\rho_{n}]^{\top}$ such that the vector values that are permuted according to $\operatorname{\bm{\rho}}$, $\textbf{{w}}_{\operatorname{\bm{\rho}}}=[w_{\rho_{1}},\dots,w_{\rho_{n}}]^{\top}$, are increasing as $w_{\rho_{1}}\leq\dots\leq w_{\rho_{n}}$. Let $\operatorname{\bm{\rho}}^{-1}$ be the inverse of permutation $\operatorname{\bm{\rho}}$, i.e., a permutation whose $\rho_{i}$-th element is $i$ for $i=1,\dots,n$.
@@
Throughout this paper, we consider rank function $r(\textbf{{w}})$ that evaluates the position of each $w_{i}$ based on sorting in an ascending order. If we need to address the ranking in descending order, we can formulate it as $r(-\textbf{{w}})$.
@@ B.1 Settings of Baselines
Regarding linear regression methods (LR-1 and LR-2), to avoid overfitting due to the large number of input features, we employed the ridge regression model in scikit-learn [Pedregosa et al., 2011]. For the meta-learner methods (i.e., the S-Learner (SL), the T-Learner (TL), the X-Learner (XL), the DR-Learner (DRL)) and the tree-based methods (CF and CF DML) we used the EconML Python package [Battocchi et al., 2019]. As the base learners for SL, TL, XL, and DRL, we chose random forest because we empirically observed that it achieved the best performance among the three model candidates: random forest, gradient boosting, and support vector machine.
We evaluated the performance of the neural network methods (TARNet and GANITE), with the existing implementations [Curth and van der Schaar, 2021, Yoon et al., 2018].  [^4] https://github.com/AliciaCurth/CATENets  [^5] https://github.com/jsyoon0823/GANITE By contrast, we implemented the DRCFR method using PyTorch [Paszke et al., 2019].
With PSW, we trained propensity score model $\pi(\textbf{{X}})$ beforehand, using a paired sample {$(a_{i},\textbf{{x}}_{i})$} $\subset\mathcal{D}$. We formulated $\pi(\textbf{{X}})$ using a three-layered FNN, as with our method and DRCFR. After computing the IPW weights based on the trained propensity score model, we performed Pareto smoothing over them with the psislw() function in a Python package called ArviZ [Kumar et al., 2019]. Using the Pareto-smoothed weights, we learned two encoders $\Delta(\textbf{{X}})$ and $\Upsilon(\textbf{{X}})$, as well as outcome prediction models $h^{0}$ and $h^{1}$.
@@ B.2 Synthetic Data
Following Hassanpour and Greiner [2020], we prepared synthetic datasets. [^6] We used their implementation in https://www.dropbox.com/sh/vrux2exqwc9uh7k/AAAR4tlJLScPlkmPruvbrTJQa?dl=0.
We first drew the values of features $\textbf{{X}}\in\operatorname{\mathbb{R}}^{d}$ from standard multivariate Gaussian distribution:
@@
Next, we sampled the values of treatment $A$ and outcome $Y$ using $\textbf{{X}}_{\Gamma}\in\operatorname{\mathbb{R}}^{d/3}$, $\textbf{{X}}_{\Delta}\in\operatorname{\mathbb{R}}^{d/3}$, and $\textbf{{X}}_{\Upsilon}\in\operatorname{\mathbb{R}}^{d/3}$, which are the feature subsets in $\textbf{{X}}=[\textbf{{X}}_{\Gamma},\textbf{{X}}_{\Delta},\textbf{{X}}_{\Upsilon}]^{\top}$. In particular, we employed $\psi=[\textbf{{X}}_{\Gamma},\textbf{{X}}_{\Delta}]^{\top}\in\operatorname{\mathbb{R}}^{2d/3}$ to generate $A$’s values as
@@
where $\textbf{{c}}_{Y^{0}}\in\operatorname{\mathbb{R}}^{2d/3}$ and $\textbf{{c}}_{Y^{1}}\in\operatorname{\mathbb{R}}^{2d/3}$ are the coefficient vectors drawn from $\mathcal{N}(0,1)$, symbol $\odot$ denotes the element-wise product (a.k.a., Hadamard product), and $\epsilon\sim\mathcal{N}(0,1)$ is scalar standard Gaussian noise. Using the values of potential outcomes $Y^{0}$ and $Y^{1}$, we computed the values of outcome $Y$ by $Y=AY^{1}+(1-A)Y^{0}$.
@@ Appendix C Additional Experimental Results
In Section 4.2, as with Hassanpour and Greiner [2020], we present the performance on relatively low-dimensional synthetic data. To further investigate its performance, we ran additional experiments using high-dimensional synthetic data.
Here we generated the synthetic data in the same way (described in Section B.2) except that the number of features $d$ was set to $d=600,1200,\dots,3000$.
Figure 4 shows the results. Again, our method achieved better CATE estimation performance than DRCFR and TARNet, thus showing that it successfully improved the estimation performance of DRCFR by effectively correcting the IPW weights by differentiable Pareto smoothing.
Regarding the feature representation learning performance (Figure 4 (a) - (c)), both DRCFR and our method outperformed TARNet. As illustrated in Figure 4 (a), however, with DRCFR and our method, the difference among the absolute parameter values in learned encoder $\Gamma(\textbf{{X}})$ decreased, as the number of features $d$ increased, indicating that distinguishing the representation of instrumental variables $\Gamma(\textbf{{X}})$ from that of the confounders $\Delta(\textbf{{X}})$ was difficult for both methods. This difficulty arises partly because binary treatment $A$ can be accurately predicted solely from the representation of the confounders (i.e., $\Delta(\textbf{{X}})$) when there are a sufficient number of confounders $\textbf{{X}}_{\Delta}$.
@@
Figure 4: Learned encoder parameter differences and test PEHEs on synthetic data: (a): value difference of $\mathbf{W}^{1}$ in encoder $\Gamma(\textbf{{X}})$; (b): value difference of $\mathbf{W}^{1}$ in encoder $\Delta(\textbf{{X}})$; (c): value difference of $\mathbf{W}^{1}$ in encoder $\Upsilon(\textbf{{X}})$; (d) test PEHEs. With TARNet, since it learns single encoder, we computed all parameter value differences with weight matrix in same encoder.